# A Unified Framework for Inference with General Missingness Patterns and Machine Learning Imputation

### v1

@@ 3.1 Method
Suppose $\mathbf{W}_{k}=\mathbf{W}$ are the same across $k$, then $\mathbf{W}^{\ast}=\left(\sum^{K}_{j=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,j}}\right)\times\left[\left(\sum^{K}_{j=1}\sum^{K}_{j^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,j},{\bm{\gamma}}_{1,j^{\prime}}}\right)+\left(\sum^{K}_{j=1}\Sigma_{{\bm{\gamma}}_{2,j}}\right)\right]^{-1}$ leads to an optimal $\Sigma_{\text{PS-PPI}}$. Accordingly, the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is
@@
The $\mathbf{W}^{\ast}$ defined in Corollary 3.2 is controlled by three components: 1) $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$: the covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and each of the $\widehat{{\bm{\gamma}}}_{1,k}$; 2) $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$: the covariance between any two of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; 3) $\Sigma_{{\bm{\gamma}}_{1,k}}$ and $\Sigma_{{\bm{\gamma}}_{2,k}}$: the variance of $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ under each missingness pattern. Specifically, a large weight matrix is assigned to all the estimation of zero $\widehat{\mathbf{\Delta}}_{k}$ when: 1) we have large covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$, which suggests that the imputations are faithfully reflect the missing true values; 2) we have smaller variance in $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$, which suggests that we can construct a precise estimation of zero with small variance; 3) we have small covariances between any pair of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, which suggests that we can efficiently aggregate across different missingness patterns without introducing significantly additional uncertainty.
@@
The theoretical results from Theorem 1 and Corollary 3.2 allow us to compute the PS-PPI estimator and the corresponding uncertainty quantification for $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in a Z-estimation framework. In particular, $\widehat{{\bm{\theta}}}_{WCC}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$ can be obtained by solving the estimating equations in Equation (3), (5), (6). For uncertainty quantification, $\widehat{\Sigma}_{{\bm{\theta}}}$, $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ can be calculated by deriving the specific closed-form given a Z-estimation framework based on results in Equation (S1.2), (S1.3), (S1.4), (S1.5), and compute their empirical estimates. We provide a more detailed instruction along with the results for linear regression and logistic regression as examples for reference, see Section S1.4 in the Supplementary Materials for details. For those Z-estimation framework without easily-computable empirical estimates of the variance-covariance matrices, one may apply the resampling strategies to address such challenge, for example, Bootstrap and delete-d Jackknife (Efron,, 1982). Under certain technical conditions, both Bootstrap and delete-$d$ Jackknife can produce a consistent variance-covariance estimation (Shao and Wu,, 1989; Shao,, 1989; Cheng and Huang,, 2010).
It is worth noting that the PS-PPI method could be implemented in a more convenient way when there exists a well-developed package for the target Z-estimation framework used for weighted complete-case analysis. Specifically, Equation (3), (5), (6) can be viewed as separate weighted complete-case analysis on different data using the same Z-estimation framework. Hence, no significant additional efforts need to be made to compute $\widehat{{\bm{\theta}}}_{WCC}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$, as well as variance-covariance matrices $\widehat{\Sigma}_{{\bm{\theta}}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ once a weighted estimation software is available. The major effort for the implementation will be, therefore, to compute $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$. As discussed above, this could be implemented by either deriving the specific closed-form based on Equation (S1.3) and (S1.4), or applying resampling strategies. Since both $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ are variance-covariance matrices only relevant to fully-observed data, the resampling will be only conducted among fully-observed data. In this paper, we take the delete-$1$ Jackknife as our strategy to estimate the covariances. See Algorithm 1 for the full algorithmic details to obtain $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and the corresponding uncertainty estimates.
@@ 3.2 Connections to the existing literature
Connections to the setting when only outcome is missing.
We show that the PS-PPI procedure subsumes the outcome-based PPI framework in the existing literature. Specifically, assume that in addition to the fully observed data subset, there is only one missingness pattern with only the outcome, say $Y_{i1}$, missing. Furthermore, assume that the underlying missing mechanism is MCAR, i.e., the propensity score model $\bm{\pi}(\cdot)$ is a constant 2-dimensional vector $(\pi_{1},\pi_{\infty})$ across all samples, where $\pi_{1}$ is the probability of missing outcome, and $\pi_{\infty}$ is the probability of being fully observed. The PS-PPI estimator under this setting become $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})$. Note that here we have the CCA estimate $\widehat{{\bm{\theta}}}_{\text{CC}}$ instead of $\widehat{{\bm{\theta}}}_{\text{WCC}}$, because the weights are the same across all samples. In  Gronsbell et al., (2024) and Chen and Chen, (2000), their proposed estimator is $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}(\widehat{{\bm{\gamma}}}_{\text{lab}}-\widehat{{\bm{\gamma}}}_{\text{all}})$. Here, $\widehat{{\bm{\theta}}}_{\text{CC}}$ follows the same definition in this paper and $\widehat{{\bm{\gamma}}}_{\text{lab}}=\widehat{{\bm{\gamma}}}_{1,1}$ with $K=1$ missingness pattern (missing outcome only); $\widehat{{\bm{\gamma}}}_{\text{all}}$ is a variant of $\widehat{{\bm{\gamma}}}_{2,1}$ in this paper, which is estimated by all the data in ${\mathcal{D}}_{\text{obs}}$, instead of only data with ${\mathcal{C}}_{i}=1$; $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter matrix estimated from the data. We show the asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{Chen}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ Proposition 2.
@@
Assume that only the outcome, say $Y_{i1}$, is possibly missing, we have that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and $\widehat{{\bm{\theta}}}_{\text{Chen}}$ are asymptotically equivalent in the linear regression and MCAR setting considered in Gronsbell et al., (2024).
Proposition 2 is proved by showing that their influence functions are the same. Note that in $\widehat{{\bm{\theta}}}_{\text{Chen}}$, $\widehat{{\bm{\gamma}}}_{\text{all}}$ is estimated using both partially and fully-observed data. This is different from $\widehat{{\bm{\gamma}}}_{2,1}$ in $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ which is estimated using only the missing-outcome subset of data.  Gronsbell et al., (2024) argue that this modification guarantees a better efficiency when $\mathbf{W}$ is the identity matrix than the original PPI procedure (Angelopoulos et al., 2023a, ). However, when an optimal $\mathbf{W}^{\ast}$ is selected, Proposition 2 indicates that such efficiency improvement is absorbed into the $\mathbf{W}^{\ast}$, thereby the two estimators become asymptotically equivalent. We elaborate this equivalence using the example of perfect predictions, i.e., the predictions are exactly the true values. Specifically, we have the following remark:
@@
Remark 3.1 can be proved by deriving the closed-form of variance-covariance matrices under linear regression and obtain the closed form of $\Sigma_{\text{PS-PPI}}$. From Remark 3.1, when perfect predictions are available, $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\gamma}}}_{\text{all}}$, which is equivalent to the results from the underlying full data since the predictions are perfect. On the other hand, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\pi}_{1}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})=\widehat{\pi}_{\infty}\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}+\widehat{\pi}_{1}\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$. Here $\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}$ is the CCA estimate obtained from the complete-case subset; $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ is the CCA estimate obtained from the underlying full data of the missing-outcome subset of data. We are able to replace $\widehat{{\bm{\gamma}}}_{2,1}$ with $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ because the perfect predictions make the imputed dataset exactly the same as the underlying full data. As a result, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be viewed as a weighted sum of the full data estimator from complete-case and missing-outcome subset of data, respectively; the weights are the proportion of the sample size of each subsets $\widehat{\pi}_{\infty}$ and $\widehat{\pi}_{1}$. Therefore, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be seen as an inverse variance weighted version of $\widehat{{\bm{\theta}}}_{\text{Chen}}$.
@@
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ also shares core insights with other existing outcome-based PPI literature (Angelopoulos et al., 2023b, ; Miao et al.,, 2023; Gan et al.,, 2024), although they may not be asymptotically equivalent. In particular, Angelopoulos et al., 2023b ; Miao et al., (2023); Gan et al., (2024) introduce tunable weights within the estimating equations, and derive similar estimators with guarantees of being more efficient than CCA. All of these methods are based on the outcome-missing setting. The proposed PS-PPI procedure introduces the tunable weights when combining the estimators of zero obtained from each missingness pattern. Each estimator of zeros can be obtained by completing two weighted CCA steps (for obtaining $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$), each of which is easy-to-implement by leveraging existing weighted CCA software, often available, for the same Z-estimation problem upon careful construction of imputed datasets following Algorithm 1.
Connections to the blockwise covariates missing setting.
@@
In addition, the PS-PPI procedure also subsumes the blockwise covariates missing scenario considered in Kluger et al., (2025). In the framework of PS-PPI, the missing blockwise covariates setting is similar to the missing outcome settings, where we only have two missingness patterns: 1) fully observed; 2) $s$ different variables in $\bm{Y}_{i}$ are missing simultaneously. Furthermore, the missing mechanism is assumed to be MAR while the propensity score model is assumed to be known. In this setting, we show the asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{PTD}}$ proposed in Kluger et al., (2025) and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in Proposition 3:
@@
In the context of semiparametric inference theory, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ belongs to the class of AIPW estimators, which can be shown by its influence functions derived from the proof of Theorem 1 (detailed in Section S1.5). In the context of semiparametric theory, one may be interested in searching for the most efficient estimators within the AIPW class estimators under specific settings (Robins et al.,, 1994, 1995; Chaudhuri and Guilkey,, 2016). However, under general missingness patterns and with an assumption of MAR, the optimal AIPW Z-estimators, despite being the most efficient estimator in integrating the predicted surrogates, suffer from difficulties in implementation as it generally does not have closed-form and would require iterative computation to solve. We refer readers to Robins et al., (1994); Tsiatis, (2006); Chaudhuri and Guilkey, (2016) for a more in-depth discussion. However, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ does not aim to guarantee optimal influence functions within the AIPW class estimators. Rather, the major advantage of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is its ability to be easily applied in analysis without significant additional efforts in the implementation. With the rapid development of deep learning and machine learning techniques, highly precise ML predictions become increasingly available. Enabling analysts to easily combine these ML-based prediction surrogates with their statistical models to perform the analysis without additional significant implementation efforts is necessary. As discussed above, our proposed PS-PPI procedure circumvents the implementation challenges by allowing researchers to apply existing statistical software designed for weighted complete-case analysis to obtain the PS-PPI estimator: each component of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ (i.e., $\widehat{{\bm{\theta}}}_{\text{WCC}},\{\widehat{{\bm{\gamma}}}_{1,k}\},\{\widehat{{\bm{\gamma}}}_{2,k}\}$) can be obtained from weighted complete-case analysis algorithms and the weights can be obtained by resampling techniques (see Algorithm 1).
@@
Multiple imputation based on Rubin’s rule is a well-established methodology that has been widely applied to account for the uncertainty introduced by imputation models to guarantee valid downstream statistical inference (Little and Rubin,, 2019). Multiple imputation often assumes a scenario where the observed data are the only resources, and imputation models are fitted based on these observed data to generate repeated imputations. To guarantee valid inference from these multiply-imputed datasets, multiple imputation imposes assumptions on the imputation models (Rubin,, 1987). Specifically, multiple imputation requires a proper imputation procedure—the unbiasedness of the point estimates and the estimated variance averaged over multiple imputed dataset, as well as the between-imputation variability—to derive a valid statistical inference (Murray,, 2018; Rubin,, 1996). Such scenario is different from the scenario assumed in PS-PPI, as well as the broader prediction-based inference literature. In additional to the observed data, current prediction-based inference literature assumes that analysts also have access to some potentially precise predictions generated from models that are trained by external data. Commonly, these external models are ML/DL-based, and therefore are either challenging to manipulate to satisfy any assumptions, or beyond analysts’ control. This is the core reason why prediction-based inference literature, including this work, do not impose any assumptions towards the prediction models. However, under the MAR missing mechanism, this weak assumption on the quality of the prediction models comes at the expense of requiring the correct specification of the propensity score model. We present a numerical experiment to investigate PS-PPI’s robustness against propensity score misspecification in Section 4 and leave detailed discussion towards propensity score model specification in Section 5.
@@
1:An observed dataset ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$ with $K$ missingness patterns (${\mathcal{C}}=1,\ldots,K$) and a subset of complete data ($\mathcal{C}$$\infty$); a set of external machine learning prediction models $h_{k}$ imputing missing entries under each missingness pattern; a propensity score models (either known or estimated from the data) $\widehat{\bm{\pi}}(\cdot)\in\mathbb{R}^{K+1}$, with each dimension $\pi_{k}(\cdot)$ representing the probability of ${\mathcal{C}}=k$. A Z-estimation procedure $f({\tt complete\leavevmode\nobreak\ data};{\tt weighting\leavevmode\nobreak\ function})$ that supports weighted complete-case analysis and returns the estimates and the variance estimates for the parameter of interest $\bm{\theta}$.

### v2

@@ 3.1 Method
Suppose $\mathbf{W}_{k}=\mathbf{W}$ across $k$, then
$\mathbf{W}^{\ast}=\left(\sum^{K}_{j=1}\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,j}}\right)\times\left[\left(\sum^{K}_{j=1}\sum^{K}_{j^{\prime}=1}\Sigma_{{\bm{\gamma}}_{1,j},{\bm{\gamma}}_{1,j^{\prime}}}\right)+\left(\sum^{K}_{j=1}\Sigma_{{\bm{\gamma}}_{2,j}}\right)\right]^{-1}$
leads to an optimal $\Sigma_{\text{PS-PPI}}$. Accordingly, the asymptotic variance of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is
@@
The $\mathbf{W}^{\ast}$ defined in Corollary 3.2 is controlled by three components: 1) $\Sigma_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$: the asymptotic covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and each of the $\widehat{{\bm{\gamma}}}_{1,k}$; 2) $\Sigma_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$: the asymptotic covariance between any two of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$; 3) $\Sigma_{{\bm{\gamma}}_{1,k}}$ and $\Sigma_{{\bm{\gamma}}_{2,k}}$: the asymptotic variance of $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$ under each missingness pattern. Specifically, a large weight matrix is assigned to $\widehat{\mathbf{\Delta}}_{k}$ when: 1) we have a large covariance between $\widehat{{\bm{\theta}}}_{\text{WCC}}$ and $\widehat{{\bm{\gamma}}}_{1,k}$, which occurs when the imputations faithfully reflect the missing true values; 2) we have small variances of $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{2,k}$, which means that we can construct a precise estimation of zero with small variances; 3) we have small covariances between any pairs of the $\widehat{{\bm{\gamma}}}_{1,k}$ and $\widehat{{\bm{\gamma}}}_{1,k^{\prime}}$, which means that we can efficiently aggregate across different missingness patterns without introducing too much extra uncertainty.
@@
We present the full algorithmic details in Algorithm 1 to obtain $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and the corresponding uncertainty estimates.
The theoretical results from Theorem 1 and Corollary 3.2 allow us to compute $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and quantify its uncertainty in a Z-estimation framework. In particular, $\widehat{{\bm{\theta}}}_{WCC}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$ can be obtained by solving the estimating equations in Equation (3), (5), (6). For uncertainty quantification, $\widehat{\Sigma}_{{\bm{\theta}}}$, $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$, $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$, and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ may be calculated by deriving closed-form expressions in a Z-estimation framework based on results in Equation (S1.3), (S1.4), (S1.5), (S1.6), and compute their empirical estimates. We provide more detailed calculations for linear regression and logistic regression as special cases in Section S1.4 in the Supplementary Materials. For general Z-estimation problems without readily computable empirical estimates of the variance-covariance matrices, resampling techniques such as Bootstrap and delete-d Jackknife (efron1982jackknife) can be used. Under certain technical conditions, both Bootstrap and delete-$d$ Jackknife can produce a consistent variance-covariance estimation (shao1989general; shao1989efficiency; cheng2010bootstrap).
It is worth noting that the PS-PPI method can be implemented provided there exists an off-the-shelf package for the target Z-estimation problem with options for weighted analyses based on complete data. Specifically, Equation (3), (5), (6) can be viewed as separate weighted complete-case analyses on different data using the same Z-estimation framework. Hence, no significant additional efforts need to be made to compute $\widehat{{\bm{\theta}}}_{\text{WCC}}$, $\widehat{{\bm{\gamma}}}_{1,k}$, $\widehat{{\bm{\gamma}}}_{2,k}$, as well as variance-covariance matrices $\widehat{\Sigma}_{{\bm{\theta}}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{2,k}}$ once a weighted estimation software is available. The major effort for the implementation will be, therefore, to compute $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$. As discussed above, this can be implemented by either deriving the specific closed-form based on Equation (S1.4) and (S1.5) in the Supplementary Materials, or applying resampling techniques. Since both $\widehat{\Sigma}_{{\bm{\theta}},{\bm{\gamma}}_{1,k}}$ and $\widehat{\Sigma}_{{\bm{\gamma}}_{1,k},{\bm{\gamma}}_{1,k^{\prime}}}$ are variance-covariance matrices only relevant to the fully-observed data, the resampling will only be conducted among the fully-observed data. In this paper, we use the delete-$1$ Jackknife to estimate the covariances.
@@ 3.2 Connections to the existing literature
Connections to the setting when only the outcome is missing.
We show that the PS-PPI procedure subsumes the outcome-based prediction-based inference framework in the existing literature. Specifically, assume that in addition to the fully observed data subset, there is only one missingness pattern with only the outcome, say $Y_{i1}$, missing. Furthermore, assume that the underlying missing mechanism is MCAR, i.e., the propensity score model $\bm{\pi}(\cdot)$ is a constant 2-dimensional vector $(\pi_{1},\pi_{\infty})$ across all the samples, where $\pi_{1}$ is the probability of missing outcome, and $\pi_{\infty}$ is the probability of being fully observed. The PS-PPI estimator under this setting become $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})$. Note that here we have the CCA estimate $\widehat{{\bm{\theta}}}_{\text{CC}}$ instead of $\widehat{{\bm{\theta}}}_{\text{WCC}}$, because the weights are the same across all samples. In gronsbell2024another and chen2000unified, their proposed estimator is $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\mathbf{W}}_{\text{Chen}}(\widehat{{\bm{\gamma}}}_{\text{lab}}-\widehat{{\bm{\gamma}}}_{\text{all}})$. Here, $\widehat{{\bm{\theta}}}_{\text{CC}}$ follows the same definition in the present paper and $\widehat{{\bm{\gamma}}}_{\text{lab}}=\widehat{{\bm{\gamma}}}_{1,1}$ with $K=1$ missingness pattern (missing outcome only); $\widehat{{\bm{\gamma}}}_{\text{all}}$ is a variant of $\widehat{{\bm{\gamma}}}_{2,1}$ in this paper, which is estimated by all the data in ${\mathcal{D}}_{\text{obs}}$, instead of only data with ${\mathcal{C}}_{i}=1$; $\widehat{\mathbf{W}}_{\text{Chen}}$ is a tuning parameter matrix estimated from the data. We show the asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{Chen}}$ and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in Proposition 2.
@@
Assume that only the outcome, say $Y_{i1}$, is possibly missing, we have that $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ and $\widehat{{\bm{\theta}}}_{\text{Chen}}$ are asymptotically equivalent in the linear regression (squared error loss) and MCAR setting considered in gronsbell2024another.
Proposition 2 is proved by showing that their influence functions are the same. Note that in $\widehat{{\bm{\theta}}}_{\text{Chen}}$, $\widehat{{\bm{\gamma}}}_{\text{all}}$ is estimated using both partially and fully-observed data. This is different from $\widehat{{\bm{\gamma}}}_{2,1}$ in $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$, which is estimated using only the missing-outcome subset of data. gronsbell2024another argue that this modification guarantees better efficiency than the original PPI procedure (angelopoulos2023prediction). However, when an optimal $\mathbf{W}^{\ast}$ is selected, Proposition 2 indicates that such efficiency improvement is absorbed into the $\mathbf{W}^{\ast}$, and thereby the two estimators become asymptotically equivalent. We elaborate this equivalence using the example of perfect predictions, i.e., the predictions are exactly the true values. Specifically, we have the following remark:
@@
Remark 3.1 can be proved by deriving the closed-form of the variance-covariance matrices under linear regression and obtaining the closed form of $\Sigma_{\text{PS-PPI}}$. From Remark 3.1, when perfect predictions are available, $\widehat{{\bm{\theta}}}_{\text{Chen}}=\widehat{{\bm{\gamma}}}_{\text{all}}$, which is equivalent to the results from the underlying full data since the predictions are perfect. On the other hand, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}=\widehat{{\bm{\theta}}}_{\text{CC}}-\widehat{\pi}_{1}(\widehat{{\bm{\gamma}}}_{1,1}-\widehat{{\bm{\gamma}}}_{2,1})=\widehat{\pi}_{\infty}\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}+\widehat{\pi}_{1}\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$. Here $\widehat{{\bm{\theta}}}^{(\infty)}_{\text{CC}}$ is the CCA estimate obtained from the complete-case subset; $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ is the CCA estimate obtained from the underlying full data of the missing-outcome subset of data. We are able to replace $\widehat{{\bm{\gamma}}}_{2,1}$ with $\widehat{{\bm{\theta}}}^{(1)}_{\text{CC}}$ because the perfect predictions make the imputed dataset exactly the same as the underlying full data. As a result, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be viewed as a weighted sum of the full data estimator from the complete-case and the missing-outcome subset of data, respectively; the weights are the proportions of each subset: $\widehat{\pi}_{\infty}$ and $\widehat{\pi}_{1}$. Therefore, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ can be viewed as an inverse variance weighted version of $\widehat{{\bm{\theta}}}_{\text{Chen}}$.
@@
$\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ also shares core insights with other existing outcome-based prediction-based inference literature (angelopoulos2023PPI++; miao2023assumption; gan2024prediction), although they may not be asymptotically equivalent. In particular, angelopoulos2023PPI++; miao2023assumption; gan2024prediction introduce tunable weights within the estimating equations, and derive similar estimators with guarantees of being more efficient than CCA. All of these methods are based on the outcome-missing setting.
Connections to missing multiple covariates setting.
@@
In addition, the PS-PPI procedure also subsumes the missing covariates scenario studied in kluger2025prediction, which considered two missingness patterns: 1) complete cases (all $p$ variables in $\bm{Y}_{i}$); 2) $s$ different variables in $\bm{Y}_{i}$ are missing simultaneously. Furthermore, the missing mechanism was assumed to be MAR with known probabilities for each missingness pattern. In the above missing covariates setting, we can establish the following asymptotic equivalence between $\widehat{{\bm{\theta}}}_{\text{PTD}}$ proposed in kluger2025prediction and $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$:
@@
In the context of semiparametric inference theory, $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ belongs to the class of AIPW estimators, which can be shown by its influence functions derived from the proof of Theorem 1 (detailed in Section S1.5 in the Supplementary Materials). In the context of semiparametric theory, one may be interested in searching for the most efficient estimators within the AIPW class of estimators under specific settings (robins1994estimation; robins1995analysis; chaudhuri2016gmm). However, under general missingness patterns and with an assumption of MAR, the optimal AIPW Z-estimators, despite being the most efficient estimator in integrating the predicted surrogates, suffer from difficulties in implementation as they generally do not have a closed-form and would require iterative computation to solve. We refer readers to robins1994estimation, tsiatis2006semiparametric, and chaudhuri2016gmm for a more in-depth discussion. The proposed $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ in this paper does not aim to guarantee optimal influence functions within the AIPW class estimators. Rather, the major advantage of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ is its ability to be easily applied in analysis without significant additional efforts in the implementation. With the rapid development of deep learning and machine learning techniques, highly precise ML predictions become increasingly available. Enabling analysts to easily combine these ML-based prediction surrogates with their statistical models to perform the analysis without additional significant implementation efforts is necessary. Our proposed PS-PPI procedure circumvents the implementation challenges by allowing researchers to apply existing statistical software designed for weighted complete-case analysis to obtain the PS-PPI estimator: each component of $\widehat{{\bm{\theta}}}_{\text{PS-PPI}}$ (i.e., $\widehat{{\bm{\theta}}}_{\text{WCC}},\{\widehat{{\bm{\gamma}}}_{1,k}\},\{\widehat{{\bm{\gamma}}}_{2,k}\}$) can be obtained from weighted complete-case analysis algorithms and the weights can be obtained by resampling techniques (see Algorithm 1).
@@
Multiple imputation based on Rubin’s rule is a well-established methodology that has been widely applied to account for the uncertainty introduced by imputation models to guarantee valid downstream statistical inference (little2019statistical). Multiple imputation often assumes a scenario where the observed data are the only resource, and imputation models are fitted based on these observed data to generate repeated imputations. To guarantee valid inference from these multiply-imputed datasets, multiple imputation imposes assumptions on the imputation models (rubin1987multiple). Such a scenario is different from the scenario assumed in PS-PPI, as well as the broader prediction-based inference literature. In addition to the observed data, current prediction-based inference literature assumes that analysts also have access to some potentially precise predictions generated from models that are trained by external data. Commonly, these external models are ML or even deep learning (DL)-based. Therefore, they are either challenging to manipulate to satisfy any assumptions or beyond analysts’ control. This is the core reason why prediction-based inference literature, including this work, does not impose any assumptions on the prediction models. However, under the MAR missing mechanism, this weak assumption on the quality of the prediction models comes at the expense of requiring the correct specification of the propensity score model. We present a numerical experiment to investigate PS-PPI’s robustness against propensity score misspecification in Section 4 and leave a detailed discussion on propensity score model specification in Section 6.
@@
1:An observed dataset ${\mathcal{D}}_{\text{obs}}=\{(G_{{\mathcal{C}}_{i}}(\bm{Y}_{i}),{\mathcal{C}}_{i})\}^{N}_{i=1}$ with $K$ missingness patterns (${\mathcal{C}}=1,\ldots,K$) and a subset of complete data ($\mathcal{C}$$\infty$); a set of external machine learning prediction models $h_{k}$ imputing missing entries under each missingness pattern; a propensity score models (either known or estimated from the data) $\widehat{\bm{\pi}}(\cdot)\in\mathbb{R}^{K+1}$, with each dimension $\pi_{k}(\cdot)$ representing the probability of ${\mathcal{C}}=k$. A Z-estimation procedure $f({\tt complete\penalty 10000\ data};{\tt weighting\penalty 10000\ function})$ that supports weighted complete-case analysis and returns the estimates and the variance estimates for the parameter of interest $\bm{\theta}$.